# The HST-Hyperion Survey: Environmental Imprints on the Stellar-Mass Function at $z\sim 2.5$

### v1

@@ 4.3 Creating and Fitting the Stellar Mass Functions
Figure 8: SMF-ratios of Hyperion Peaks and Outskirts – The SMFs of the Hyperion outskirts (magenta) and peaks (orange) after being normalized by the coeval field, and linear fits to both sets assuming $\log_{10}(\phi/\phi_{\rm field})=\beta\log_{10}(M_{*})+\log_{10}(A)$. We additionally show results from the Elentàri proto-supercluster at $z\sim 3.3$ (Forrest et al. 2024b). In this framework, a flat line corresponds to an SMF in an overdense region with an identical shape to the coeval field, whereas a positive slope corresponds to an overdense region with a larger fraction of higher- to lower-mass galaxies.
@@ 5.1 Trends Within the Hyperion SMF
Qualitatively, the shape of the total SMF is interesting in that we see a slight dip at $\log_{10}(M_{*}/M_{\odot})\sim 10.5$, and it is thus well-fit by a double-Schechter function with a characteristic stellar mass at the point of the dip. However, the double-Schechter function has two additional degrees of freedom, and therefore has a slightly higher reduced-$\chi^{2}$ value and slightly lower BIC (Table 3). As such, it is somewhat ambiguous as to if the total SMF of Hyperion is more consistent with a single- or double-Schechter fit.
@@ 5.2 Broader Context of the Hyperion SMF
The results of our SMF analysis suggest galaxies in the most overdense peaks of Hyperion are undergoing – or have undergone – some process which enhances the buildup of stellar mass relative to field galaxies. This trend of generally increasing stellar mass in the most overdense peaks is consistent with results of similar studies at intermediate redshifts ($0.6<z<1.5$; Tomczak et al. 2017; van der Burg et al. 2020), as well as other studies of protostructures at higher redshifts ($z>2$; Shimakawa et al. 2018a, b). Additionally, in Figure 8 we make a more direct comparison with results from Forrest et al. (2024b), who used a similar VMC method to study the Elentári proto-supercluster at $z\sim 3.3$ in the COSMOS field. While a direct comparison between similar overdensities at different redshifts is difficult due to cosmic variance (see Figure 3), the SMFs of both Hyperion and Elentári suggest a similar trend of increasing stellar mass with increasing overdensity in high redshift proto-superclusters.
@@
These trends likely reflect a broader shift in how environment influences galaxy evolution at high redshifts. It is well established that galaxies in clusters at low-to-intermediate redshifts ($z\lesssim 1.5$) tend to be quiescent and gas-poor, having built up their stellar mass at some earlier time (Gómez et al. 2003; von der Linden et al. 2010; Muzzin et al. 2012; Tomczak et al. 2019; Old et al. 2020). However, recent studies have shown that protostructures at higher redshifts exhibit a fundamentally different relationship to the star formation rate (SFR) of their constituent galaxies. For instance, Lemaux et al. (2022) find a weak but statistically significant positive correlation between SFR and overdensity for $2\leq z\leq 5$, suggesting that early dense environments may in fact stimulate star formation, rather than suppress it. Likewise, a stacking analysis of infrared observations of protostructures revealed enhanced SFR densities (SFRDs) for protostructures in the redshift range $2\leq z\leq 4$ (Kubo et al. 2019; Popescu et al. 2023). In a more dedicated spectroscopic follow-up of the Taralay protocluster at $z\sim 4.57$, Staab et al. (2024) find an enhanced SFRD which they attribute to both a higher density of star forming galaxies and an elevated SFR in said galaxies.
@@
The results of our analysis in Hyperion are consistent with – and are perhaps a direct consequence of – this positive correlation between overdensity and SFRs at higher redshifts. While it makes sense that elevated levels of star formation in high redshift protostructures lead to more massive galaxies, the full process by which this might happen remains somewhat unclear. However, we offer a brief outline using results of recent studies through which an enhancement of SFR in high redshift overdense environments may result in the SMFs observed in Hyperion.
@@
By definition, overdense environments at high redshifts ($z\gtrsim 4$) have a higher number density of galaxies than the coeval field. Dynamically, galaxies in these protostructure likely have slightly elevated velocity dispersions than the field, as a result of the larger gravitational potential of the overdense environment. It is therefore fair to assume that these galaxies experience more frequent galaxy-galaxy interactions than their counterparts in the field. At lower redshifts ($z\lesssim 1$), it has been shown this “galaxy harassment” can excite transient periods of star formation in dusty galaxies (Moore et al. 1996, 1998). Although in a recent study, Shah et al. (2022) show the typical SFR-enhancement in interacting galaxies is lower at higher redshifts ($1<z<3$) than at lower redshifts, even a modest increase in SFR produced by continual close encounters could result in a moderately elevated stellar mass in the $\sim 1.5$ Gyr between $z\sim 5$ and $z\sim 2.5$.
@@
We now turn back to $z\sim 2.5$ to see if this picture holds for Hyperion and other protostructures. Within Hyperion, Giddings et al. (2025) find that the fraction of galaxies with close kinematic companions is nearly double that of the field, meaning galaxy-galaxy interactions appear to be elevated. Additionally, Gururajan et al. (2025) find a tentative positive correlation between SFR and overdensity within Hyperion, inline with the trend reported in Lemaux et al. (2022), and which may be expected if there are more frequent close encounters. Interestingly, they also find that the gas fraction ($M_{\rm gas}/M_{*}$) of member galaxies within Hyperion tends to decrease with increasing overdensity, and appears uncorrelated in the most overdense peaks. Taken together with the increased SFRs, this tentatively implies an enhanced star formation efficiency ($\rm SFE\equiv SFR/M_{gas}$) within Hyperion, which is conceivably driven by close encounters of member galaxies. The flattening of gas-fraction trends in the most overdense peaks of Hyperion, combined with the decreasing star formation seen in other protostructure peaks at $z\lesssim 2.5$ (Shimakawa et al. 2018a, b), could signal the onset of a hot intracluster medium (ICM) in this redshift range. A recent study of the Spiderweb protocluster at $z\sim 2.16$ indeed detects signatures of an ICM via the thermal Sunyaev–Zeldovich effect (Sunyaev & Zeldovich 1972; Di Mascolo et al. 2023). However, despite existing X-ray observations of Hyperion, it remains unclear whether an ICM has yet begun to develop (Wang et al. 2016; Champagne et al. 2021).
@@
Correspondingly, the overdense peaks of higher redshift structures, such as Hyperion, may be ideal locations for major galaxy mergers. As discussed previously, Giddings et al. (2025) find an elevated number of close kinematic pairs within Hyperion and estimate an average merging timescale of $\sim 3-10$ Gyr. This timescale is consistent with the $\sim 3$ Gyr separating Hyperion at $z\sim 2.5$ and the structures observed in Tomczak et al. (2017) at $z\sim 1$. Additionally, elevated fractions of major mergers in overdense protostructures are consistent with observations suggesting an increased AGN fraction at the redshift of Hyperion (Shah et al. 2024b).
Likely, the local environment of a galaxy in the high redshift universe mediates enhanced stellar mass growth in more than one way. While the SMF of Hyperion presented in this paper provides compelling evidence that the environment does play a role at some level in promoting accelerated mass growth in the early universe, it is not enough to make a definitive argument as to what method dominates the process. Given the two mechanisms presented here – enhanced SFRs via galaxy harassment and major mergers – we might expect to find different quiescent fractions at different redshifts depending on which mechanism is primarily driving galaxy evolution. Whereas close encounters and harassment may lead to a more gradual quenching as the gas reservoir is heated and exhausted, major mergers may drive more rapid quenching through stellar feedback from starbursts and AGN feedback.
Given the idea that overdense environments in the early universe can promote accelerated stellar mass growth, looking at quiescent fractions in and around these environments may provide insight into underlying physical mechanisms. Lin et al. (2010b) find little correlation between local environment and quiescent fraction over $1.5<z<2.5$ in the COSMOS field, while a more recent study by Edward et al. (2024) finds a relative excess of low-mass ($\log_{10}(M_{*}/M_{\odot})<10$) quiescent galaxies in protostructures at $2<z<2.5$, including Hyperion. However, this analysis relies solely on photometric classifications and does not explore the trend as a function of overdensity. On the contrary, results from the ZFOURGE survey suggest the quiescent fraction is higher in more overdense environments up to $z\sim 2$ for galaxies with $\log_{10}(M_{*}/M_{\odot})>9.5$ (Kawinwanichakij et al. 2017), and up to $z\sim 2.4$ for galaxies with $\log_{10}(M_{*}/M_{\odot})>10.2$ (Hartzenberg et al. 2023). At higher redshifts still, Forrest et al. (2024a) find a quiescent fraction consistent with that of the field within the Elentári proto-supercluster, though dense regions of the structure show evidence of enhanced quiescent fractions – especially at higher stellar masses (McConachie et al. 2022, 2025).
@@ 6 Conclusion
In this work, we construct the galaxy stellar mass function (SMF) in the Hyperion proto-supercluster at $z\sim 2.5$ by leveraging a combination of data from the COSMOS2020 catalog, various ground-based spectroscopic surveys, and a new targeted HST-grism survey. To our knowledge, this is the most complete SMF in a single proto-supercluster at $z\gtrsim 2$ to date. The additional spectroscopy and improved photometric redshifts enable us to re-populate the existing Voronoi Monte Carlo (VMC) overdensity map of Hyperion with a much larger, more secure galaxy sample. This enriched map lets us compare, for the first time, the SMFs of Hyperion’s overdense peaks and its lower-density outskirts (Figure 7).
@@
    The overdense peaks in Hyperion contain a markedly larger fraction of high-mass galaxies than the coeval field (Figure 8). Specifically, we find the peaks have number densities of massive galaxies ($M_{*}\sim 10^{11}\,\,M_{\odot}$) which are $\sim 10\times$ higher than the field, compared to only $\sim 3\times$ higher for lower mass galaxies ($M_{*}\sim 10^{9.5}\,\,M_{\odot}$).
@@ Acknowledgements.
DS is grateful for support through Prof. Colby Haggerty’s NSF-CAREER grant AGS-2338131, which enabled the completion of this slight scientific detour. BF acknowledges support from JWST-GO-02913.001-A. DCB is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303800. DCB is also supported by the UC Chancellor’s Postdoctoral Fellowship. LG acknowledge the FONDECYT regular project number 1230591, the ANID - MILENIO - NCN2024_112, and the ANID BASAL project FB210003 for financial support. Much of this work is based on photometry from the COSMOS2020 catalog, and we thus thank the COSMOS team for their efforts. Some of the material presented in this paper is based upon work supported by the National Science Foundation under Grant No. 1908422. This work is also based on observations collected at the European Southern Observatory under ESO programmes 175.A0839, 179.A-2005, and 185.A-0791. This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. Supported by the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America. This work is based on observations obtained with MegaPrime/MegaCam, a joint project of CFHT and CEA/IRFU, at the Canada-France-Hawaii Telescope (CFHT) which is operated by the National Research Council (NRC) of Canada, the Institut National des Science de l’Univers of the Centre National de la Recherche Scientifique (CNRS) of France, and the University of Hawaii. This work is based in part on data products produced at Terapix available at the Canadian Astronomy Data Centre as part of the Canada-France-Hawaii Telescope Legacy Survey, a collaborative project of NRC and CNRS. This work is based, in part, on observations made with the Spitzer Space Telescope, which is operated by the Jet Propulsion Laboratory, California Institute of Technology under a contract with NASA. This research is based in part on data collected at the Subaru Telescope, which is operated by the National Astronomical Observatory of Japan. Some of the data presented herein were obtained at Keck Observatory, which is a private 501(c)3 non-profit organization operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Native Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain.
Software: Astropy (Astropy Collaboration et al. 2022), Matplotlib (Hunter 2007), NumPy (Harris et al. 2020), SciPy (Virtanen et al. 2020), statsmodels (Seabold & Perktold 2010).

### v2

@@ 4.3 Creating and fitting the stellar mass functions
Figure 9: The SMFs of the Hyperion outskirts (magenta) and peaks (orange) after being normalized by the coeval field, and linear fits to both sets assuming $\log_{10}(\phi/\phi_{\rm field})=\beta\log_{10}(M_{*})+\log_{10}(A)$. We additionally show the results from the Elentàri proto-supercluster at $z\sim 3.3$ (Forrest et al. 2024b). In this framework, a flat line corresponds to an SMF in an overdense region with an identical shape to the coeval field, whereas a positive slope corresponds to an overdense region with a larger fraction of higher- to lower-mass galaxies.
@@ 5.1 Trends within the Hyperion SMF
Qualitatively, the shape of the total SMF is interesting in that we see a slight dip at $\log_{10}(M_{*}/M_{\odot})\sim 10.5$, and it is thus well fit by a double-Schechter function with a characteristic stellar mass at the point of the dip. However, the double-Schechter function has two additional degrees of freedom, and therefore has a slightly higher reduced-$\chi^{2}$ value and slightly lower BIC (Table 3). As such, it is somewhat ambiguous as to if the total SMF of Hyperion is more consistent with a single- or double-Schechter fit.
@@ 5.2 Broader context of the Hyperion SMF
The results of our SMF analysis suggest galaxies in the most overdense peaks of Hyperion are undergoing – or have undergone – some process which enhances the buildup of stellar mass relative to field galaxies. This trend of generally increasing stellar mass in the most overdense peaks is consistent with results of similar studies at intermediate redshifts ($0.6<z<1.5$; Tomczak et al. 2017; van der Burg et al. 2020), as well as other studies of protostructures at higher redshifts ($z>2$; Shimakawa et al. 2018a, b). Additionally, in Figure 9 we make a more direct comparison with results from Forrest et al. (2024b), who used a similar VMC method to study the Elentári proto-supercluster at $z\sim 3.3$ in the COSMOS field. While a direct comparison between similar overdensities at different redshifts is difficult due to cosmic variance (see Figure 4), the SMFs of both Hyperion and Elentári suggest a similar trend of increasing stellar mass with increasing overdensity in high-redshift proto-superclusters.
@@
These trends likely reflect a broader shift in how environment influences galaxy evolution at high redshifts. It is well established that galaxies in clusters at low to intermediate redshifts ($z\lesssim 1.5$) tend to be quiescent and gas-poor, having built up their stellar mass at some earlier time (Gómez et al. 2003; von der Linden et al. 2010; Muzzin et al. 2012; Tomczak et al. 2019; Old et al. 2020). However, recent studies have shown that protostructures at higher redshifts exhibit a fundamentally different relationship to the star formation rate (SFR) of their constituent galaxies. For instance, Lemaux et al. (2022) find a weak but statistically significant positive correlation between SFR and overdensity for $2\leq z\leq 5$, suggesting that early dense environments may in fact stimulate star formation, rather than suppress it. Likewise, a stacking analysis of infrared observations of protostructures revealed enhanced SFR densities (SFRDs) for protostructures in the redshift range $2\leq z\leq 4$ (Kubo et al. 2019; Popescu et al. 2023). In a more dedicated spectroscopic follow-up of the Taralay protocluster at $z\sim 4.57$, Staab et al. (2024) find an enhanced SFRD which they attribute to both a higher density of star-forming galaxies and an elevated SFR in said galaxies.
@@
The results of our analysis in Hyperion are consistent with – and are perhaps a direct consequence of – this positive correlation between overdensity and SFRs at higher redshifts. While it makes sense that elevated levels of star formation in high-redshift protostructures lead to more massive galaxies, the full process by which this might happen remains somewhat unclear. However, we offer a brief outline using results of recent studies through which an enhancement of SFR in high-redshift overdense environments may result in the SMFs observed in Hyperion.
@@
By definition, overdense environments at high redshifts ($z\gtrsim 4$) have a higher number density of galaxies than the coeval field. Dynamically, galaxies in these protostructure likely have slightly elevated velocity dispersions than the field, as a result of the larger gravitational potential of the overdense environment. It is therefore fair to assume that these galaxies experience more frequent galaxy-galaxy interactions than their counterparts in the field. At lower redshifts ($z\lesssim 1$), it has been shown this galaxy harassment can excite transient periods of star formation in dusty galaxies (Moore et al. 1996, 1998). Although in a recent study, Shah et al. (2022) show the typical SFR-enhancement in interacting galaxies is lower at higher redshifts ($1<z<3$) than at lower redshifts, even a modest increase in SFR produced by continual close encounters could result in a moderately elevated stellar mass in the $\sim 1.5$ Gyr between $z\sim 5$ and $z\sim 2.5$.
@@
We now turn back to $z\sim 2.5$ to see if this picture holds for Hyperion and other protostructures. Within Hyperion, Giddings et al. (2026) find that the fraction of galaxies with close kinematic companions is nearly double that of the field, meaning galaxy-galaxy interactions appear to be elevated. Additionally, Gururajan et al. (2025) find a tentative positive correlation between SFR and overdensity within Hyperion, inline with the trend reported in Lemaux et al. (2022), and which may be expected if there are more frequent close encounters. Interestingly, they also find that the gas fraction ($M_{\rm gas}/M_{*}$) of member galaxies within Hyperion tends to decrease with increasing overdensity, and appears uncorrelated in the most overdense peaks. Taken together with the increased SFRs, this tentatively implies an enhanced star formation efficiency ($\rm SFE\equiv SFR/M_{gas}$) within Hyperion, which is conceivably driven by close encounters of member galaxies. The flattening of gas-fraction trends in the most overdense peaks of Hyperion, combined with the decreasing star formation seen in other protostructure peaks at $z\lesssim 2.5$ (Shimakawa et al. 2018a, b), could signal the onset of a hot intracluster medium (ICM) in this redshift range. A recent study of the Spiderweb protocluster at $z\sim 2.16$ indeed detects signatures of an ICM via the thermal Sunyaev–Zeldovich effect (Sunyaev and Zeldovich 1972; Di Mascolo et al. 2023). However, despite existing X-ray observations of Hyperion, it remains unclear whether an ICM has yet begun to develop (Wang et al. 2016; Champagne et al. 2021).
@@
Correspondingly, the overdense peaks of higher-redshift structures, such as Hyperion, may be ideal locations for major galaxy mergers. As discussed previously, Giddings et al. (2026) find an elevated number of close kinematic pairs within Hyperion and estimate an average merging timescale of $\sim 3-10$ Gyr. This timescale is consistent with the $\sim 3$ Gyr separating Hyperion at $z\sim 2.5$ and the structures observed in Tomczak et al. (2017) at $z\sim 1$. Additionally, elevated fractions of major mergers in overdense protostructures are consistent with observations suggesting an increased AGN fraction at the redshift of Hyperion (Shah et al. 2025).
Likely, the local environment of a galaxy in the high-redshift universe mediates enhanced stellar mass growth in more than one way. While the SMF of Hyperion presented in this paper provides compelling evidence that the environment does play a role at some level in promoting accelerated mass growth in the early universe, it is not enough to make a definitive argument as to what method dominates the process. Given the two mechanisms presented here – enhanced SFRs via galaxy harassment and major mergers – we might expect to find different quiescent fractions at different redshifts depending on which mechanism is primarily driving galaxy evolution. Whereas close encounters and harassment may lead to a more gradual quenching as the gas reservoir is heated and exhausted, major mergers may drive more rapid quenching through stellar feedback from starbursts and AGN feedback.
Given the idea that overdense environments in the early universe may promote accelerated stellar mass growth, looking at quiescent fractions in and around these environments might provide insight into underlying physical mechanisms. Lin et al. (2010b) find little correlation between local environment and quiescent fraction over $1.5<z<2.5$ in the COSMOS field, while a more recent study by Edward et al. (2024) finds a relative excess of low-mass ($\log_{10}(M_{*}/M_{\odot})<10$) quiescent galaxies in protostructures at $2<z<2.5$, including Hyperion. However, this analysis relies solely on photometric classifications and does not explore the trend as a function of overdensity. On the contrary, results from the ZFOURGE survey suggest the quiescent fraction is higher in more overdense environments up to $z\sim 2$ for galaxies with $\log_{10}(M_{*}/M_{\odot})>9.5$ (Kawinwanichakij et al. 2017), and up to $z\sim 2.4$ for galaxies with $\log_{10}(M_{*}/M_{\odot})>10.2$ (Hartzenberg et al. 2023). At higher redshifts still, Forrest et al. (2024a) find a quiescent fraction consistent with that of the field within the Elentári proto-supercluster, though dense regions of the structure show evidence of enhanced quiescent fractions – especially at higher stellar masses (McConachie et al. 2022, 2025).
@@ 6 Conclusion
In this work, we construct the galaxy stellar mass function (SMF) in the Hyperion proto-supercluster at $z\sim 2.5$ by leveraging a combination of data from the COSMOS2020 catalog, various ground-based spectroscopic surveys, and a new targeted HST-grism survey. To our knowledge, this is the most complete SMF in a single proto-supercluster at $z\gtrsim 2$ to date. The additional spectroscopy and improved photometric redshifts enable us to repopulate the existing Voronoi Monte Carlo (VMC) overdensity map of Hyperion with a much larger, more secure galaxy sample. This enriched map lets us compare, for the first time, the SMFs of Hyperion’s overdense peaks and its lower-density outskirts (Figure 8).
@@
    The overdense peaks in Hyperion contain a markedly larger fraction of high-mass galaxies than the coeval field (Figure 9). Specifically, we find that the peaks have number densities of massive galaxies ($M_{*}\sim 10^{11}\,\,M_{\odot}$) that are $\sim 10\times$ higher than the field, compared to only $\sim 3\times$ higher for lower-mass galaxies ($M_{*}\sim 10^{9.5}\,\,M_{\odot}$).
@@ Acknowledgements.
We would like to thank the anonymous referee for their helpful comments. DS is grateful for support through Prof. Colby Haggerty’s NSF-CAREER grant AGS-2338131, which enabled the completion of this slight scientific detour. BF acknowledges support from JWST-GO-02913.001-A. DCB is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303800. DCB is also supported by the UC Chancellor’s Postdoctoral Fellowship. LG acknowledge the FONDECYT regular project number 1230591, the ANID - MILENIO - NCN2024_112, and the ANID BASAL project FB210003 for financial support. Much of this work is based on photometry from the COSMOS2020 catalog, and we thus thank the COSMOS team for their efforts. Some of the material presented in this paper is based upon work supported by the National Science Foundation under Grant No. 1908422. This work is also based on observations collected at the European Southern Observatory under ESO programmes 175.A0839, 179.A-2005, and 185.A-0791. This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. Supported by the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America. This work is based on observations obtained with MegaPrime/MegaCam, a joint project of CFHT and CEA/IRFU, at the Canada-France-Hawaii Telescope (CFHT) which is operated by the National Research Council (NRC) of Canada, the Institut National des Science de l’Univers of the Centre National de la Recherche Scientifique (CNRS) of France, and the University of Hawaii. This work is based in part on data products produced at Terapix available at the Canadian Astronomy Data Centre as part of the Canada-France-Hawaii Telescope Legacy Survey, a collaborative project of NRC and CNRS. This work is based, in part, on observations made with the Spitzer Space Telescope, which is operated by the Jet Propulsion Laboratory, California Institute of Technology under a contract with NASA. This research is based in part on data collected at the Subaru Telescope, which is operated by the National Astronomical Observatory of Japan. Some of the data presented herein were obtained at Keck Observatory, which is a private 501(c)3 non-profit organization operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Native Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain.
Software: Astropy (astropy), Matplotlib (MatPlotLib), NumPy (Numpy), SciPy (Scipy), statsmodels (statsmodels).